# Generalised height pairings and the Albanese kernel

### v1

@@ 2.3. Examples
    $\mathcal{C}$ is the category of finite dimensional continuous $\mathbb{Q}_{p}$-representations of the Galois group of a finite extension of $\mathbb{Q}_{\ell}$, $\ell\neq p$. We will denote this by $M_{g}(V_{0},V_{1},V_{2})$.
@@
    $\mathcal{C}$ is the category of finite dimensional de Rham representation of the Galois group of a finite extension of $\mathbb{Q}_{p}$. We will denote this by $M_{g}(V_{0},V_{1},V_{2})$.
@@
    $\mathcal{C}$ is the category of filtered $\phi$-modules over a finite extension of $\mathbb{Q}_{p}$.
@@
This is also a consequence of the weight monodromy conjecture for abelian varieties (see [Nek93] or [BD21, Lemma 3.4]. ∎
@@
$1\to H^{1}_{g}(K_{v},W)\to M_{g}(\mathbb{Q}_{p},V,W)\to H^{1}_{g}(K_{v},V)\times H^{1}_{g}(K_{v},V^{*}\otimes W)\to 1$
@@
[…]
@@ 2.5. Generalised heights
In [BD21], it was observed that this construction admits a natural generalisation, relavent to quadratic Chabauty. Let $W$ be a Galois stable quotient 555later, we will also consider general quotients of $V^{\otimes 2}$. By Lemma 8, the difference is not significant for nonabelian Chabauty. of $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{et}}(X,\mathbb{Q}_{p})\stackrel{{\scriptstyle\cup}}{{\longrightarrow}}H^{2}_{\mathrm{et}}(X,\mathbb{Q}_{p}))^{*}$. The local generalised height is a function
@@
For the purposes of this paper, the choice of local generalised heights will not be important.
@@
The map $h$ is bilinear in the following sense. Extend $h$ to $\mathbb{Q}_{p}[M_{g}(\mathbb{Q}_{p},V,W)]$ linearly, i.e.
@@
A conjecture of Bloch and Kato [BK90] implies that the map
$(\pi_{1*},\pi_{2*}):M_{g}(\mathbb{Q}_{p},V,W)\to H^{1}_{g}(K,V)\times H^{1}_{g}(K,V^{*}\otimes W)$
is surjective. This would imply that the generalised height in fact defines a pairing
@@
which factors through $H^{1}_{f}(K,V)\otimes_{R}H^{1}_{f}(K,V^{*}\otimes W)$, where $R$ acts on $H^{1}_{f}(K,V^{*}\otimes W)\simeq\operatorname{Ext}^{1}_{f}(V,W)$ via its action on $V$. A particularly interesting case is that of mixed extension with graded pieces $\mathbb{Q}_{p},V^{\oplus n},W$, for $W$ some quotient of $V^{\otimes 2}$. Then we have an obvious inclusion of $\mathrm{Mat}_{n}(\mathbb{Q}_{p})$ in $\mathrm{End}_{G_{K}}(V)$, and an isomorphism
@@
$0\to\frac{\oplus_{v}H^{1}(K_{v},W)}{\operatorname{loc}H^{1}(K,W)}\to H^{2}_{c}(K,W)\to\mathrm{Ker}(H^{2}(K,W)\to\oplus_{v}H^{2}(K_{v},W))\to 0$
@@ 2.6. The mixed extensions associated to fundamental groups
$\mathbb{Z}[\pi_{1}(X(\mathbb{C}),b)]/I^{n+1}\otimes_{\mathbb{Z}}\mathbb{Q}_{\ell},$
@@
Let $E_{n}(X;b,z)$ denote the fibre of $E_{n}(X;b)$ at $z$. By the above, an element of $E_{n}(X;b,z)$ can be thought of as a $\mathbb{Q}_{\ell}$-linear combination of homotopy classes of paths from $b$ to $z$ (although this perspective will not be useful in what follows). $E_{n}(X;b,z)$ inherits an action of the Galois group of $K$. By the above, $E_{n}(X,b)$ admits a filtration corresponding to the filtration by powers $I^{k}$ of the augmentation ideal on the universal enveloping algebra. We denote by $I^{k}E_{n}(X;b,z)$ the fibre at $z$ of the local system $I^{k}E_{n}(X;b)$. Where there is no possibility of confusion we write the fibre at $b$ as simply $E_{n}(X;b),I^{k}/I^{j}(b)$, etc. When $n=2$, we will drop the subscript $n$ and denote this simply as $E(X;b),E(X;b,z)$, etc.
@@
Let $i_{1}$ and $i_{2}$ be the inclusions $V\hookrightarrow V^{*}\otimes(V^{\otimes 2}/\mathbb{Q}_{p}(1))$. Then we have the following identities in $\mathrm{Ext}^{1}_{\operatorname{Gal}(K)}(V,V^{\otimes 2}/\mathbb{Q}_{p}(1))$.
@@
If $X$ is hyperelliptic, and $P$ is a Weierstrass point, then $[IE(P,P)]=0$ in $\mathrm{Ext}^{1}_{\operatorname{Gal}(K)}(V,V^{\otimes 2}/\mathbb{Q}_{p}(1))$.
@@
$[IE(P,P)]=w^{*}[IE(P,P)]$
On the other hand $w^{*}$ acts as $-1$ on $V$ and as $1$ on $V^{\otimes 2}/\mathbb{Q}_{p}(1)$, so we deduce that $02[IE(P,P)]=$ 0. ∎
From this we deduce the following property of $E(b,z)$.
@@
If $X$ is a hyperelliptic curve, then $E(b,z)$ is an extension of $[z-b]\in H^{1}_{f}(\mathbb{Q},V)$ by $i_{1*}(z-W)+i_{2*}(W-b)$ in $H^{1}_{f}(\mathbb{Q},V^{\otimes 3}(-1)/V)$, where $i_{1}$ and $i_{2}$ are the maps
@@
Lemma 9.
The generalised height function satisfies Galois descent, i.e. if $L|K$ is a finite Galois extension of number fields the diagram
@@
commutes.
In this way, we can think of the generalised height as defining a function
$\mathrm{colim}_{K|\mathbb{Q}}\mathrm{Im}(\mathbb{Q}_{p}[H^{1}_{g}(U(\mathbb{Q}_{p},V,W)]\to H^{1}_{g}((H^{1}_{f}(K,V)\otimes H^{1}_{f}(K,V^{*}\otimes W))))^{\operatorname{Gal}(K|\mathbb{Q})}\to H^{1}_{f}(\mathbb{Q},W),$
where the colimit is over finite Galois extensions $K|\mathbb{Q}$. We now consider how to define height functions on divisors, which should be viewed as generalisations of the usual quadratic $p$-adic height function on an elliptic curve.
@@ 2.7. Mixed extensions from curves in $X\times X$
Fix a closed immersion $(f_{1},f_{2}):C\hookrightarrow X\times X$. We now describe two ways to construct mixed extensions in $M(\mathbb{Q}_{p},V,V^{\otimes 2})$ from such curves. Suppose
$(f_{1}^{*},f_{2}^{*}):H^{0}(X,\Omega)^{2}\to H^{0}(C,\Omega)$
is surjective. Then the Jacobian of $C$ is isogenous to $J\times J\times A$ for some abelian variety $A$. Hence its $\mathbb{Q}_{p}$-Tate module is isomorphic to $V^{\oplus 2}\oplus V_{A}$. We obtain a surjection
$\wedge^{2}T_{p}\mathrm{Jac}(C)\to V^{\otimes 2}.$
Hence for any pair of points $x,y\in C(K)$, we obtain a mixed extension $E_{C}(x,y)\in M(\mathbb{Q}_{p},V^{\oplus 2}\oplus V_{A},V^{\otimes 2})$. The associated generalised height defines a map which factors through
$H^{1}_{f}(\mathbb{Q},V)\otimes H^{1}_{f}(\mathbb{Q},V^{\otimes 3}(-1))\oplus H^{1}_{f}(\mathbb{Q},V_{A})\otimes H^{1}_{f}(\mathbb{Q},V_{A}^{*}\otimes V^{\otimes 2}).$
@@
The class of $E_{C}(x,y)$ in $H^{1}_{f}(\mathbb{Q},V_{A}^{*}\otimes V^{\otimes 2})$ is independent of $x$ and $y$.
@@
Hence we obtain a map
$\mathrm{PDiv}^{0}(C)\to H^{1}_{f}(K,V)\otimes H^{1}_{f}(K,V^{\otimes 3}(-1)).$
@@
Let $C\subset X\times X$ be a geometrically connected curve and $D=\sum n_{i}(P_{i},Q_{i})$ a principal divisor on $C$. Then the image of $\sum n_{i}P_{i}Q_{i}\in J(\mathbb{Q})^{\otimes 2}$ in $H^{1}(\mathbb{Q},V)\otimes H^{1}(\mathbb{Q},V^{*}\otimes V^{\otimes 2}/\mathbb{Q}_{p}(1))$ lies in the image of $H^{1}_{f}(\mathbb{Q},U(\mathbb{Q}_{p},V^{\oplus n},V^{\otimes 2}/\mathbb{Q}_{p}(1)))$ for some $n$.
@@
Note that the cohomological interpretation of the generalised heights constructed from the curves $C$ is to look at the Galois representations associated to appropriate subquotients of $H^{2}(X^{2}-C\cup\{x_{i}\}\times X\cup X\times\{y_{j}\})$. This is one motivation for the constructions in the next section.

### v2

@@ 2.3. Examples
    $\mathcal{C}$ is the category of finite dimensional continuous $\mathbb{Q}_{p}$-representations of the Galois group of a finite extension of $\mathbb{Q}_{\ell}$, $\ell\neq p$. We will denote this by $M_{g}(V_{0},V_{1},V_{2})$, or sometimes $M_{g,v}(V_{0},V_{1},V_{2})$.
@@
    $\mathcal{C}$ is the category of finite dimensional de Rham, or crystalline, representation of the Galois group of a finite extension of $\mathbb{Q}_{p}$. We will denote these categories by $M_{g}(V_{0},V_{1},V_{2})$ and $M_{f}(V_{0},V_{1},V_{2})$ respectively.
@@
    $\mathcal{C}$ is the category of filtered $\phi$-modules over a finite extension of $\mathbb{Q}_{p}$. We will denote this category by $M(V_{0},V_{1},V_{2})$.
@@
This is also a consequence of the weight monodromy conjecture for abelian varieties (see [Nek93] or [BD21, Lemma 3.4]). ∎
@@
$1\to H^{1}_{g}(K_{v},W)\to M_{g}(\mathbb{Q}_{p},V,W)\to H^{1}_{g}(K_{v},V)\times H^{1}_{g}(K_{v},V^{*}\otimes W)\to 1.$
@@
In the case where $v$ is prime to $p$, this is proved in [BD21] in the case of crystalline representations. We recall the argument: one shows that there is a commutative diagram with exact rows whose vertical maps are all isomorphisms
$H^{1}_{f}(K_{v},W)$$H^{1}_{f}(K_{v},U(\mathbb{Q}_{p},V,W))$$H^{1}_{f}(K_{v},V)\times H^{1}_{f}(K_{v},V^{*}\otimes W)$$\operatorname{Ext}^{1}_{f}(\mathbb{Q}_{p},W)$$M_{f}(\mathbb{Q}_{p},V,W)$$\operatorname{Ext}^{1}_{f}(\mathbb{Q}_{p},V)\times\operatorname{Ext}_{f}(V,W)$$\operatorname{Ext}^{1}(\mathbb{Q}_{p},D(W))$$M(\mathbb{Q}_{p},D(V),D(W))$$\operatorname{Ext}^{1}(\mathbb{Q}_{p},D(V))\times\operatorname{Ext}(D(V),D(W))$$D_{\operatorname{dR}}(W)/F^{0}$$F^{0}\backslash U(K_{v},D_{\operatorname{dR}}(V),D_{\operatorname{dR}}(W))$$D_{\operatorname{dR}}(V)/F^{0}\times D_{\operatorname{dR}}(V^{*}\otimes W)/F^{0}$}
where $D(.):=D_{\mathrm{cris}}(.)$, the ext groups and mixed extensions in the third row are in the category of filtered $\phi$-modules over $K_{v}$, the middle vertical maps are given by the $D_{\mathrm{cris}}$ functors, and the bottom vertical maps are given by Kim’s classification of unipotent torsors in the category of filtered $\phi$-modules (see [BD21, §3.3]). Since the bottom extends to a short exact sequence, so does the top.
@@ 2.5. Generalised heights
In [BD21], it was observed that this construction admits a natural generalisation, relevant to quadratic Chabauty. Let $W$ be a Galois stable quotient 555later, we will also consider general quotients of $V^{\otimes 2}$. By Lemma 8, the difference is not significant for nonabelian Chabauty. of $\mathrm{Ker}(\wedge^{2}H^{1}_{\mathrm{et}}(X,\mathbb{Q}_{p})\stackrel{{\scriptstyle\cup}}{{\longrightarrow}}H^{2}_{\mathrm{et}}(X,\mathbb{Q}_{p}))^{*}$. The local generalised height is a function
@@
In [BD21], a recipe for constructing a local height at a prime $v$ above $p$ depending on a choice of a splitting of the Hodge filtration on $H^{1}_{\operatorname{dR}}(X_{K_{v}}/K_{v})$. For the purposes of this paper, the choice of local generalised heights will not be important.
@@
It follows from the definitions that the map $h$ is bilinear in the following sense [BD21, Lemma 3.8]. Extend $h$ to $\mathbb{Q}_{p}[M_{g}(\mathbb{Q}_{p},V,W)]$ linearly, i.e.
@@
A conjecture of Bloch and Kato [BK90, Conjecture 5.3 (i)] implies that the map
(5) $(\pi_{1*},\pi_{2*}):M_{g}(\mathbb{Q}_{p},V,W)\to H^{1}_{g}(K,V)\times H^{1}_{g}(K,V^{*}\otimes W)$
is surjective. Indeed, this conjecture implies that the localisation map
$H^{1}(G_{K,T},W^{*}(1))\to\oplus_{v\in T}H^{1}(G_{K_{v}},W^{*}(1))$
is injective, (see e.g. [BD21, Lemma 2.4]) and by Poitou–Tate duality this implies that the localisation map
$H^{1}(G_{K,T},W^{*}(1))\to\oplus_{v\in T}H^{2}(G_{K_{v}},W)$
is injective. Surjectivity of (5) then follows from local surjectivity, which is proved in Lemma 4.
Hence conjecturally the generalised height in fact defines a pairing
@@
which factors through $H^{1}_{f}(K,V)\otimes_{R}H^{1}_{f}(K,V^{*}\otimes W)$, where $R$ acts on $H^{1}_{f}(K,V^{*}\otimes W)\simeq\operatorname{Ext}^{1}_{f}(V,W)$ via its action on $V$. As is shown in loc. cit., a generalised height will be $R$-equivariant if it is constructed with respect to $R$-equivariant splittings of the Hodge filtration. A particularly interesting case is that of mixed extensions with graded pieces $\mathbb{Q}_{p},V^{\oplus n},W$, for $W$ some quotient of $V^{\otimes 2}$. Then we have an obvious inclusion of $\mathrm{Mat}_{n}(\mathbb{Q}_{p})$ in $\mathrm{End}_{G_{K}}(V)$, and an isomorphism
@@
$0\to\frac{\oplus_{v}H^{1}(K_{v},W)}{\operatorname{loc}H^{1}(K,W)}\to H^{2}_{c}(K,W)\to\mathrm{Ker}(H^{2}(K,W)\to\oplus_{v}H^{2}(K_{v},W))\to 0.$
@@ 2.6. The mixed extensions associated to fundamental groups
$\mathbb{Z}[\pi_{1}(X(\mathbb{C}),b)]/I^{n+1}\otimes_{\mathbb{Z}}\mathbb{Q}_{p},$
@@
Let $E_{n}(X;b,z)$ denote the fibre of $E_{n}(X;b)$ at $z$. By the above, an element of $E_{n}(X;b,z)$ can be thought of as a $\mathbb{Q}_{p}$-linear combination of homotopy classes of paths from $b$ to $z$ (although this perspective will not be used in what follows). $E_{n}(X;b,z)$ inherits an action of the Galois group of $K$. By the above, $E_{n}(X,b)$ admits a filtration corresponding to the filtration by powers $I^{k}$ of the augmentation ideal on the universal enveloping algebra. We denote by $I^{k}E_{n}(X;b,z)$ the fibre at $z$ of the local system $I^{k}E_{n}(X;b)$. Where there is no possibility of confusion we write the fibre at $b$ as simply $E_{n}(X;b),I^{k}/I^{j}(b)$, etc. When $n=2$, we will drop the subscript $n$ and denote this simply as $E(X;b),E(X;b,z)$, etc.
@@
Let $i_{1}$ and $i_{2}$ be the inclusions $V\hookrightarrow V^{*}\otimes(V^{\otimes 2}/\mathbb{Q}_{p}(1))$ corresponding to left and right multiplication. Then we have the following identities in $\mathrm{Ext}^{1}_{\operatorname{Gal}(K)}(V,V^{\otimes 2}/\mathbb{Q}_{p}(1))$.
@@
If $X$ is hyperelliptic, and $P$ is a Weierstrass point, then $[IE(P)]=0$ in $\mathrm{Ext}^{1}_{\operatorname{Gal}(K)}(V,V^{\otimes 2}/\mathbb{Q}_{p}(1))$.
@@
$IE(P)]=w^{*}[IE(P)]$
On the other hand $w^{*}$ acts as 1-1 on $V$ and as $1$ on $V^{\otimes 2}/\mathbb{Q}_{p}(1)$, so we deduce that $2[IE(P)]=0$. ∎
@@
If $X$ is a hyperelliptic curve, then $E(b,z)$ is an extension of $z-b]\in H^{1}_{f}(\mathbb{Q},V)$ by $i_{1*}(z-W)+i_{2*}(W-b)$ in $H^{1}_{f}(\mathbb{Q},V^{\otimes 3}(-1)/V)$, where $i_{1}$ and $i_{2}$ are the maps
@@
Proof.
This follows from Lemma 5 and Lemma 6. ∎
@@
It follows from the definitions that the generalised height function satisfies Galois descent, i.e. if $L|K$ is a finite Galois extension of number fields the diagram
@@
commutes. In this way, we can think of the generalised height as defining a function
$\mathrm{colim}_{K|\mathbb{Q}}\mathrm{Im}\left(\mathbb{Q}_{p}[H^{1}_{g}(U(\mathbb{Q}_{p},V,W)]^{\operatorname{Gal}(K|\mathbb{Q})}\to(H^{1}_{f}(K,V)\otimes H^{1}_{f}(K,V^{*}\otimes W))^{\operatorname{Gal}(K|\mathbb{Q})}\right)$
$\to(\oplus_{v}H^{1}_{g}(\mathbb{Q}_{v},W)/H^{1}_{g}(\mathbb{Q},W),$
where the colimit is over finite Galois extensions $K|\mathbb{Q}$. We now consider how to define height functions on divisors, which could be viewed as generalisations of the usual quadratic $p$-adic height function on an elliptic curve.
@@ 2.7. Mixed extensions from curves in $X\times X$
Fix a closed immersion $(f_{1},f_{2}):C\hookrightarrow X\times X$, where $C$ is a geometrically irreducible curve. We obtain a direct sum decomposition
$V_{C}\simeq V_{A}\oplus V_{B}$
where $V_{C}$ is the $\mathbb{Q}_{p}$-Tate module of the Jacobian of $C$, $V_{A}$ is the $\mathbb{Q}_{p}$-Tate module of the connected component of the identity $A$ of
$\mathrm{Ker}(\mathrm{Jac}(C)\to\mathrm{Jac}(X)^{2})$
induced by $(f_{1},f_{2})$, and $V_{B}$ is the $\mathbb{Q}_{p}$-Tate module of $\mathrm{Jac}(C)/A$. We obtain a surjection
$\wedge^{2}V_{C}\to\wedge^{2}V.$
induced by the two maps from $V_{C}$ to $V$. Hence for any pair of points $x,y\in C(K)$, we obtain a mixed extension $E_{C}(x,y)\in M(\mathbb{Q}_{p},V_{C},W)$ where $W:=\wedge^{2}V/\mathrm{NS}(\mathrm{Jac}(X)_{\overline{\mathbb{Q}}})^{*}\otimes\mathbb{Q}_{p}(1)$. Now let $R:=\mathrm{End}(B)\otimes\mathbb{Q}_{p}\times\mathrm{End}(A)\otimes\mathbb{Q}_{p}\subset\mathrm{End}_{\operatorname{Gal}(K)}(V_{C})$, let $R_{0}:=\mathrm{End}(\mathrm{Jac}(X))\otimes\mathbb{Q}_{p}$ and $R_{1}:=\mathrm{End}(A)\otimes\mathbb{Q}_{p}$. Then any associated $R$-equivariant generalised height defines a map which factors through
$H^{1}_{f}(\mathbb{Q},V_{C})\otimes_{R}H^{1}_{f}(\mathbb{Q},(V_{C})^{*}\otimes W)$
$\simeq$ $H^{1}_{f}(\mathbb{Q},V)\otimes_{R_{0}}H^{1}_{f}(\mathbb{Q},V\otimes W)\oplus H^{1}_{f}(\mathbb{Q},V_{A})\otimes_{R_{1}}H^{1}_{f}(\mathbb{Q},V_{A}^{*}\otimes W).$
@@
The class of $E_{C}(x,y)$ in $H^{1}_{f}(\mathbb{Q},V_{A}^{*}\otimes W)$ is independent of $x$ and $y$.
@@
Hence the map
$\mathbb{Q}_{p}[C(\overline{\mathbb{Q}})\times C(\overline{\mathbb{Q}})]^{\operatorname{Gal}(\overline{\mathbb{Q}}|\mathbb{Q})}\to\mathrm{colim}_{K|\mathbb{Q}}(H^{1}_{f}(K,V_{C})\otimes_{R}H^{1}_{f}(K,V_{C}^{*}\otimes W))^{\operatorname{Gal}(K|\mathbb{Q})}$
induces a map
$\theta:\mathrm{PDiv}^{0}(C)\to\mathrm{colim}_{K|\mathbb{Q}}(H^{1}_{f}(K,V)\otimes H^{1}_{f}(K,V^{*}\otimes W))^{\operatorname{Gal}(K|\mathbb{Q})}.$
@@
Let $C\subset X\times X$ be a geometrically irreducible curve, $W$ as above, and $D=\sum n_{i}(P_{i},Q_{i})$ a principal divisor on $C$. Let $h_{C}$ denote the generalised height on $H^{1}_{g}(\mathbb{Q}_{p},V_{C},W)$ with respect to an $R$-equivariant splitting $s$ of the Hodge filtration as above. Let $h$ denote the generalised height on $X$ with respect to a splitting compatible with $s$. Suppose the class of $\sum n_{i}P_{i}\otimes Q_{i}$ in $H^{1}_{f}(K,V)^{\otimes 2}\subset H^{1}_{f}(K,V)\otimes H^{1}_{f}(K,V^{*}\otimes W)$ is in the image of $\mathrm{colim}\mathbb{Q}_{p}[H^{1}_{g}(K,U(\mathbb{Q}_{p},V,W))]^{\operatorname{Gal}(K|\mathbb{Q})}$. Then
$h_{C}(D)=h(\sum n_{i}P_{i}\otimes Q_{i})$
where the height on the right hand side is the generalised height on the image of
$\mathrm{colim}\mathbb{Q}_{p}[H^{1}_{g}(K,U(\mathbb{Q}_{p},V,W))]^{\operatorname{Gal}(K|\mathbb{Q})}\to\mathrm{colim}\left(H^{1}_{f}(K,V)\otimes H^{1}_{f}(K,V^{*}\otimes W)\right)^{\operatorname{Gal}(K|\mathbb{Q})}.$
Proof.
For all primes $v$, if $M$ is a mixed extension in $H^{1}_{g}(K_{v},U(\mathbb{Q}_{p},V,W))$ then by construction we have
$h_{v}(M)=h_{C,v}(M\oplus T)$
where $T:=\mathrm{Ker}(V_{C}\stackrel{{\scriptstyle f_{1*}}}{{\longrightarrow}}V)$, and $M\oplus T$ is viewed as an element of $H^{1}_{g}(K_{v},U(\mathbb{Q}_{p},V_{C},W))$ in the obvious way.
On the other hand, we have a commutative diagram
$\mathbb{Q}_{p}[H^{1}_{g}(K,U(\mathbb{Q}_{p},V,W))]$$\mathbb{Q}_{p}[H^{1}_{g}(K,U(\mathbb{Q}_{p},V_{C},W))]$$H^{1}_{f}(K,V)\otimes_{R_{0}}H^{1}_{f}(K,V^{*}\otimes W)$$H^{1}_{f}(K,V)\otimes_{R_{0}}H^{1}_{f}(K,(V^{*}\otimes W)\oplus H^{1}_{f}(K,V_{A})\otimes_{R_{1}}H^{1}_{f}(K,V_{A}^{*}\otimes W).$}
Hence the lemma follows from Lemma 10. ∎
@@
Note that the cohomological interpretation of the generalised heights constructed from the curves $C$ is to look at the Galois representations associated to appropriate subquotients of $H^{2}(X^{2}-C\cup\{P_{i}\}\times X\cup X\times\{Q_{j}\})$. This is one motivation for the constructions in the next section.